OpenStudy (anonymous):

What is the solution of the quadratic inequality x2 − 4x − 32 ≥ 0?

OpenStudy (anonymous):

x^2-(8-4)x-32>=0 x^2-8x+4x-32>=0 x(x-8)+4(x-8)>=0 (x+4)(x-8)>=0

OpenStudy (anonymous):

now,either (x+4)>=0 x>=-4 or, x-8>=0 x>=8

OpenStudy (juanpablojr):

I think its x>=-4. but Gather more Data I might be wrong.